A coil of wire is made of 520 turns and has an area of 0.30 m^2. It is placed in a uniform magnetic field of magnitude 0.40 T. This magnetic field flips its orientation from 0.40 T in the +y-direction to 0.40 T in the −y-direction in 2.0 s. How much power is generated in the coil if it has a resistance of 1200 Ω ?


1
Expert's answer
2021-05-16T17:53:36-0400

Number of turns, "n=520"

Area of of cross section, "A=0.30\space m^2"

Emf induced, "|\epsilon|=n\dfrac{|\Delta\Phi|}{|\Delta t|}=n\dfrac{|BA-(-BA)|}{2}"

"|\epsilon|=520\times\dfrac{|2BA|}{2}=520\times0.40=208\space V"


Resistance, "R=1200\space\Omega"

Power generated, "P=\dfrac{V^2}{R}=\dfrac{(208)^2}{1200}=36.05\space W"